OpenStudy (anonymous):

Use 15◦= 45◦− 30◦to find the exact value of sin15 , cos15 , tan15

OpenStudy (nenadmatematika):

sin(x-y)=sinxcosy-sinycosx cos(x-y)=cosxcosy+sinxsiny.......tan(x-y)=sin(x-y)/cos(x-y)....try it !!!!! :D

OpenStudy (anonymous):

others are similar damn i made a mistake hold on

OpenStudy (anonymous):

\[\sin(15)=\sin(45-30)=\sin(45)\cos(30)-\sin(30)\sin(45)\] \[=\frac{\sqrt{2}}{2}\times \frac{\sqrt{3}}{2}-\frac{\sqrt{2}}{2}\times \frac{1}{2}\] \[\frac{\sqrt{6}}{4}-\frac{\sqrt{2}}{4}=\frac{\sqrt{6}-\sqrt{2}}{4}\] that's better

OpenStudy (anonymous):

what formula do i use for cos though ?

OpenStudy (nenadmatematika):

you have a formula above...just plug in 45 for x and 15 for y....try to learn!!!! :D